OpenStudy (anonymous):

Scores on a math midterm are distrubted under a standard normal curve. A randomly selected student had a score of 55, which represented a z score of -2.35. If the mean for the exam was 76, determine the standard deviation for the exam

OpenStudy (kropot72):

\[z=\frac{X-\mu}{\sigma}\] \[-2.35=\frac{55-76}{\sigma}\] \[\sigma=\frac{55-76}{-2.35}=?\] Can you calculate the value of sigma (the standard deviation) now?

OpenStudy (kropot72):

\[\frac{55-76}{-2.35}=\frac{-21}{-2.35}=\frac{21}{2.35}=?\]

OpenStudy (kropot72):

@Murry Are you there? Do you have a calculator?

OpenStudy (anonymous):

Yes i have a calculator